Consider the parametric curve given by x(t) = t^2 − 2, y(t) = 2t^3 + 3, What is the length of the arc from the point (−1, 1) to the point (2, −13)
Answer #1

# Arc dength = = I t (dy, 2 dx. GE x = 2 2 - 0 9 = 2+² +3 - 1 t² = (2+0) ( od t= (2+2 0/2 Pect the value of t in Ecs - (11) Yplease do like if this helped you